Calculate the future value of R10 000 invested for 5 years at an interest rate of 7 %, interest compounded semi-annually.... a) R14 026 b)R14 106 c) R14 900 d) R16 234
1
Expert's answer
2020-01-21T07:36:09-0500

"A= P(1+r\/n) ^ t.n"

"=10,000(1+0.07\/2)^5.2\n\n=14,105.9\n\n=R 14,106"